OpenStudy (anonymous):

What is the geometric mean of 8 and 12

OpenStudy (anonymous):

\[ \sqrt{ 8 (12) } =\sqrt{96}= 4 \sqrt 6 \]

OpenStudy (mathmale):

xoxo: I'd strongly recommend that you look up and explain to yourself on paper what "geometric mean" denotes (means). I don't know that myself, so my very first move would have been to look up "geometric mean" on the 'Net.

OpenStudy (♪chibiterasu):

Multiply the two numbers together and square root the product.

OpenStudy (mathstudent55):

The geometric mean of and b is x if \(\dfrac{a}{x} = \dfrac{x}{b} \) Solving for x, you get: \x^2 = ab\) \(x = \sqrt{ab} \)